# Upstream history quantification and scale-decomposed energy analysis for weak-to-strong adverse-pressure-gradient turbulent boundary layers

### v1

@@ 2.2 Case Selection and Analysis Methodology
Figure 3: Small and large scale contributions to velocity variances in Case 2. DNS22 contributions are not plotted due to unavailability of energy spectra data at a reasonable "Quasi-matched" location. Decomposition into small and large scales is done based on the $\lambda_{z,c}^{+}=300$ cutoff documented in Deshpande et al. (2023). All quantities are scaled in viscous units - $U_{\tau}$ and $l_{\tau}$. (a,e) Small scale contributions to $\overline{u^{2}}$ variances, (b,f) Large scale contributions to $\overline{u^{2}}$ variances, (c,g) Small scale contributions to $\overline{v^{2}}$ variances, (d,h) Large scale contributions to $\overline{v^{2}}$ variances. Colors and Line styles correspond to those in Tables 1 and 2.
@@ 3 Results
We now analyze the five matched cases described in the previous section to understand the upstream PG history and local disequilibration effects. The primary goals are to: a) establish the effect of $d\beta_{C}/dRe_{\tau}$ on the flow statistics and b) to determine whether the accumulated PG parameters (defined in equations 1, 2 and 3) capture the upstream PG history effects across a broad range of pressure gradient strengths. The success of the latter will facilitate establishing the effects of PG history on the small- and large-scale energies across weak-to-strong APG TBLs.
@@ 3.1 Low-to-moderately strong APG TBLs
We begin with Case 1 comparing the TBL datasets b1.4 and b1.0, analyzed previously by Pozuelo et al. (2022), to revisit the pre-established ability of the $\overline{\beta}_{C,\theta}$ parameter to capture the effects of APG history for moderate values of $\beta_{C}$ (Vinuesa et al., 2017). Figure 2(top) depicts the inner-scaled mean velocity profiles along with the inner-scaled $\overline{u^{2}}$ and $\overline{v^{2}}$ variances for the two data sets at Exactly-Matched Case 1. As already discussed by Pozuelo et al. (2022), the behavior of the boundary layer statistics and the greater turbulence energisation observed in the b1.0 simulation are logically accounted for by $\overline{\beta}_{C,\theta}$, consistent with the original description in Bobke et al. (2017). The mean velocity profiles reveal two key observations that illustrate the effects of upstream history. First, b1.0 having $\overline{\beta}_{C,\theta}$ of 0.84, which is more than twice that of the b1.4 (0.39), is characterized by a greater deviation from the ZPG in the log region, as seen in the inset of Figure 2(a). Similar deviations from the classical log law with increasing APG strength have been reported across low-to-high Reynolds-number TBLs in the past (Nagib and Chauhan, 2008; Monty et al., 2011; Indinger et al., 2006; Knopp et al., 2021; Pozuelo et al., 2022; Deshpande et al., 2023; Gungor et al., 2024a). Second, in the outer region, the b1.0 TBL exhibits a larger mean velocity deficit as well as higher Reynolds normal stress levels (Figures 2b and 2c), further highlighting the stronger cumulative effect of the APG associated with stronger history. Here, for brevity, we have limited our analysis only to $\overline{u^{2}}$ and $\overline{v^{2}}$ but similar trends are noted for $\overline{w^{2}}$ and $\overline{uv}$. Local disequilibration appears to play a minor role in Case 1 (relative to accumulated PG history), as the negative ${\rm d}\beta_{C}/{\rm d}Re_{\tau}$ observed for flow b1.0 would typically indicate a reduction in the APG effect (Gungor et al., 2024b).
@@
While Case 1 limits itself to a comparison of two low-$\beta$ TBLs at a matched $\beta_{C}$ and $Re_{\tau}$, in Case 2 we compare three TBLs at matched $\beta_{C}$ and $Re_{\theta}$ – one non-equilibrium TBL DNS22, in its region of rapidly increasing PG and two low-$\beta$ TBLs b1.4 and b1.0. The comparison in Figure 2 (bottom row) between the b1.0 TBL ($\overline{\beta}_{C,\theta}=0.82$) and the b1.4 TBL ($\overline{\beta}_{C,\theta}=0.50$) is consistent with the conclusion drawn from Case 1 regarding the influence of accumulated pressure gradient on TBL statistics, even when matching $Re_{\theta}$ instead of $Re_{\tau}$. We observe a greater mean velocity deficit, greater deviation from ZPG in the log region, and higher Reynolds normal stress levels in the outer region of the b1.0 TBL compared to the b1.4 TBL. Figure 3 presents scale-decomposed fluctuations for Quasi-matched values associated with Case 2 (Table 2) to analyze the small- and large-scale contributions to $\overline{u^{2}}$ and $\overline{v^{2}}$ variances decomposed based on the $\lambda^{+}_{z,c}=300$ cutoff. Here, we only plot scale-decomposed variances for b1.0 and b1.4 due to the unavailability of a reasonable "Quasi-matched" location for DNS22. Note that the slight mismatch in $\beta$ and $Re$ values for Quasi-matched cases does not influence the statistical trends, and can be confirmed by referring to Appendix B. Observing $\overline{u^{2}_{SS}}$ and $\overline{u^{2}_{LS}}$ reveals a striking difference between small- and large-scale energisation in response to increasing accumulated PG ($\overline{\beta}_{C,\theta}$). Small-scale contributions to $\overline{u^{2}}$ variances (see Figure 3a) exhibit a collapse in the inner region but a clear energisation in the outer region, aligning with the increasing $\overline{\beta}_{C,\theta}$ value. On the other hand, the large-scale contributions ($\overline{u^{2}_{LS}}$) plotted in Figure 3(b) unveil an energisation of large-scale variance contributions with increasing APG history across the entire TBL. Scale-decomposed contributions of $\overline{v^{2}}$ variances also demonstrate an energisation with $\overline{\beta}_{C,\theta}$ beyond the near-wall region, where most of the $\overline{v^{2}}$ energy is concentrated (Figures 3c,d).
@@
Figure 5: Small and large scale contributions to velocity variances in Cases 3 (top), 4 (middle), and 5 (bottom). Decomposition into small and large scales is done based on the $\lambda_{z,c}/\delta=0.5$ cutoff discussed in §$\S$2.2. All quantities are scaled with respect to outer units, $U_{e}$ and $\delta$. (a,e,i) Small scale contributions to $\overline{u^{2}}$ variances, (b,f,j) Large scale contributions to $\overline{u^{2}}$ variances, (c,g,k) Small scale contributions to $\overline{v^{2}}$ variances, (d,h,l) Large scale contributions to $\overline{v^{2}}$ variances. Colors and Line styles correspond to those in Tables 1 and 2.
@@ 3.2 TBLs imposed with strong adverse pressure gradients
For Case 3, $Re_{\theta}$ and $\beta_{ZS}$ are matched, but the higher value of $\overline{\beta}_{ZS,\theta}$ in DNS22 suggests that it experiences a more pronounced historical effect of the APG than b1.4. Indeed, figure 4(a) shows the mean velocity defect increasing with $\overline{\beta}_{ZS,\theta}$, indicating a greater cumulative influence of the APG, consistent with our observations for $\overline{\beta}_{C,\theta}$ in Figures 2(top). Similarly, Figures 4(b) and 4(c) respectively show increasing outer peaks in the $\overline{u^{2}}$ and $\overline{v^{2}}$ variances with an increase in the $\overline{\beta}_{ZS,\theta}$ parameter. It is important to note here that the choice of an outer scaling-based PG parameter restricts our analysis solely to the outer region of the boundary layer, preventing reliable inferences about the inner region. We note that although the accumulated PG parameter $\overline{\beta}_{ZS,\theta}$ qualitatively aligns with the differences in the statistics, and is consistent with previously established trends of $\overline{\beta}_{C,\theta}$, the effectiveness of using $\overline{\beta}_{ZS,\theta}$ together with matched values of $Re_{\theta}$ and $\beta_{ZS}$ cannot be conclusively determined based on this case alone.
@@
In terms of influence of local disequilibration, one can argue these effects are likely not to influence the statistics in Case 5 due to ${\rm d}{\beta}/{\rm d}{Re}$ values being of the same order of magnitude for the associated data sets (refer Table 2). For Cases 3 and 4, however, local disequilibration may play a more significant role, given that the ${\rm d}{\beta}/{\rm d}{Re}$ values differ by at least an order of magnitude between the two data sets being compared. Strong negative ${\rm d}{\beta_{ZS}}/{\rm d}{Re}_{\theta}$ for DNS22 (Case 3) and DNS16 (Case 4) imply a decreasing impact of the APG (Gungor et al., 2024b), which acts in opposition to the strong positive value of $\overline{\beta}_{ZS,\theta}$. The results for Cases 3 and 4 in Figure 4(a-f) indicate, however, that the accumulated PG effect dominates relative to the local disequilibration.
@@
With the efficacy of $\overline{\beta}_{ZS,\theta}$ and $\overline{\beta}_{ZS,ZS}$ to quantify history effects demonstrated, next we leverage this new capability to understand the historical PG trends in scale-decomposed contributions to velocity variances for these strong APG TBLs. Figure 5 presents scale-decomposed contributions to $\overline{u^{2}}$ and $\overline{v^{2}}$ variances for Cases 3 to 5. We reiterate here that outer scaling limits our analysis of flow physics solely to the outer region. The plots reveal outer region energisation of both small- and large-scale contributions of $\overline{u^{2}}$ and $\overline{v^{2}}$ with increasing $\overline{\beta}_{ZS}$ (2 and 3), across all three high-defect Cases 3-5, consistent with Case 2 (Figure 2 (below)). Further, for cases 3 and 4, the trend associated with the local disequilibration ($d{\beta}/d{Re}$; for both small and large-scales) is also consistent with that noted for Case 2 in the outer region.

### v2

@@ 2.2 Case Selection and Analysis Methodology
Availability of spectral energy data at the "Quasi-Matched" cases plays a crucial role in this study, enabling analysis of small- and large-scale energy contributions as a function of upstream PG history. Although the $\beta$ and $Re$ values do not match exactly for these comparisons, the slight mismatch does not influence the statistical trends inferred in subsequent sections. We utilize the cutoff based on the viscous-scaled spanwise wavelength ($\lambda^{+}_{z,c}=300$) proposed by Deshpande et al. (2023) to decompose the total spectral energy at each wall-normal location into small ($\overline{u^{2}_{SS}}$; $\lambda^{+}_{z,c} \leq$ 300) and large scale contributions ($\overline{u^{2}_{LS}}$; $\lambda^{+}_{z,c}$ >> 300), such that $\overline{u^{2}}$($y^{+}$) = $\overline{u^{2}_{SS}}$($y^{+}$) + $\overline{u^{2}_{LS}}$($y^{+}$), and $\overline{v^{2}}$($y^{+}$) = $\overline{v^{2}_{SS}}$($y^{+}$) + $\overline{v^{2}_{LS}}$($y^{+}$). Deshpande et al. (2023) arrived at this cutoff value iteratively, based on the well-accepted argument of viscous-scaled small-scale energy ($\overline{u^{2}_{SS}}^{+}$) exhibiting $Re_{\tau}$-invariance for $\lambda_{z,c}^{+}\leq 300$ (for the case of canonical flows; refer to figure 12 of Deshpande et al., 2023). Since this spanwise wavelength cutoff, $\lambda^{+}_{z,c}$ is constructed based on a viscous-scaled parameter that breaks down for strong APG TBLs, it becomes ineffective in higher-defect cases. This limitation is overcome by utilizing an outer-scaled spanwise wavelength ($\lambda_{z,c}/\delta$), the cutoff value for which is estimated iteratively such that the small- and large-scale velocity variance profiles nominally match those computed based on the established value of $\lambda^{+}_{z,c}$ = 300, for the low-defect Cases 1 and 2. The cutoff value of $\lambda_{z,c}/\delta$ = 0.5 shows reasonable agreement between the scale-decomposed variances, as demonstrated in the plots provided in Appendix A.
@@ 3 Results
We now analyze the five matched cases and the special Case DISEQ described in the previous section to understand the upstream PG history and local disequilibration effects. The primary goals are to: a) establish the effect of $d\beta_{C}/dRe_{\tau}$ on the flow statistics and b) to determine whether the accumulated PG parameters (defined in equations 1, 2 and 3) capture the upstream PG history effects across a broad range of pressure gradient strengths. The success of the latter will facilitate establishing the effects of PG history on the small- and large-scale energies across weak-to-strong APG TBLs.
@@ 3.1 Low-to-moderately strong APG TBLs
We begin with Case 1 comparing the TBL datasets b1.4 and b1.0, analyzed previously by Pozuelo et al. (2022), to revisit the pre-established ability of the $\overline{\beta}_{C,\theta}$ parameter to capture the effects of APG history for moderate values of $\beta_{C}$ (Vinuesa et al., 2017). Figure 2(top) depicts the inner-scaled mean velocity profiles along with the inner-scaled $\overline{u^{2}}$ and $\overline{v^{2}}$ variances for the two data sets at Exactly-Matched Case 1. As already discussed by Pozuelo et al. (2022), the behavior of the boundary layer statistics and the greater turbulence energisation observed in the b1.0 simulation are logically accounted for by $\overline{\beta}_{C,\theta}$, consistent with the original description in Bobke et al. (2017). The mean velocity profiles reveal two key observations that illustrate the effects of upstream history. First, b1.0 having $\overline{\beta}_{C,\theta}$ of 0.83, which is more than twice that of the b1.4 (0.48), is characterized by a greater deviation from the ZPG in the log region, as seen in the inset of Figure 2(a). Similar deviations from the classical log law with increasing APG strength have been reported across low-to-high Reynolds-number TBLs in the past (Nagib and Chauhan, 2008; Monty et al., 2011; Indinger et al., 2006; Knopp et al., 2021; Pozuelo et al., 2022; Deshpande et al., 2023; Gungor et al., 2024a). Second, in the outer region, the b1.0 TBL exhibits a larger mean velocity deficit as well as higher Reynolds normal stress levels (Figures 2b and 2c), further highlighting the stronger cumulative effect of the APG associated with stronger history. Here, for brevity, we have limited our analysis only to $\overline{u^{2}}$ and $\overline{v^{2}}$ but similar trends are noted for $\overline{w^{2}}$ and $\overline{uv}$. Local disequilibration appears to play a minor role in Case 1 (relative to accumulated PG history), as the negative ${\rm d}\beta_{C}/{\rm d}Re_{\tau}$ observed for flow b1.0 would typically indicate a reduction in the APG effect (Gungor et al., 2024b).
Figure 3: Premultiplied spanwise power-spectral density of the (a) streamwise, (b) wall normal velocity fluctuations ($k_{z}\phi_{uu}^{+}$ and $k_{z}\phi_{vv}^{+}$ respectively), small and large-scale contributions to velocity variances in Case 2 scaled in inner units. DNS22 contributions are not plotted due to the unavailability of power-spectral density data at a reasonable "Quasi-matched" location. Decomposition into small and large scales is done based on the $\lambda_{z,c}^{+}=300$ cutoff documented in Deshpande et al. (2023). All quantities are scaled in viscous units - $U_{\tau}$ and $l_{\tau}$. Figure also presents (c) Small scale contributions to $\overline{u^{2}}$ variances, (d) Large scale contributions to $\overline{u^{2}}$ variances, (e) Small scale contributions to $\overline{v^{2}}$ variances, (f) Large scale contributions to $\overline{v^{2}}$ variances. Colors and Line styles correspond to those in Tables 1 and 2.
@@
While Case 1 limits itself to a comparison of two low-$\beta$ TBLs at a matched $\beta_{C}$ and $Re_{\tau}$, in Case 2 we compare three TBLs at matched $\beta_{C}$ and $Re_{\theta}$ – one non-equilibrium TBL DNS22, in its region of rapidly increasing PG and two low-$\beta$ TBLs b1.4 and b1.0. The comparison in Figure 2 (bottom row) between the b1.0 TBL ($\overline{\beta}_{C,\theta}=0.82$) and the b1.4 TBL ($\overline{\beta}_{C,\theta}=0.50$) is consistent with the conclusion drawn from Case 1 regarding the influence of accumulated pressure gradient on TBL statistics, even when matching $Re_{\theta}$ instead of $Re_{\tau}$. We observe a greater mean velocity deficit, greater deviation from ZPG in the log region, and higher Reynolds normal stress levels in the outer region of the b1.0 TBL compared to the b1.4 TBL. Figure 3 presents premultiplied spanwise power-spectral density and scale-decomposed fluctuations scaled in inner units for Quasi-matched values associated with Case 2 (Table 2) to analyze the response of the turbulent scales to the changing PG conditions. The small- and large-scale contributions to $\overline{u^{2}}$ and $\overline{v^{2}}$ variances are decomposed based on the $\lambda^{+}_{z,c}=300$ cutoff. Here, we only plot the power-spectral density and scale-decomposed variances for b1.0 and b1.4 due to the unavailability of a reasonable "Quasi-matched" location for DNS22. Note that the slight mismatch in $\beta$ and $Re$ values for Quasi-matched cases does not influence the statistical trends, and can be confirmed by referring to Appendix B. Figure 3(a) shows the development of an outer peak in the $\overline{u^{2}}$ spectra of the b1.0 TBL, which is consistent with its greater $\overline{\beta}_{C,\theta}$ value, indicating a more pronounced APG effect. Furthermore, the energy in the outer region of the b1.0 spectra (both $\overline{u^{2}}$ and $\overline{v^{2}}$ energy) extends to both smaller and larger scales, further supporting this.
Now, observing $\overline{u^{2}_{SS}}$ and $\overline{u^{2}_{LS}}$ reveals a striking difference between small- and large-scale energisation in response to increasing accumulated PG ($\overline{\beta}_{C,\theta}$). Small-scale contributions to $\overline{u^{2}}$ variances (see Figure 3a) exhibit a collapse in the inner region but a clear energisation in the outer region, aligning with the increasing $\overline{\beta}_{C,\theta}$ value. On the other hand, the large-scale contributions ($\overline{u^{2}_{LS}}$) plotted in Figure 3(b) unveil an energisation of large-scale variance contributions with increasing APG history across the entire TBL. Scale-decomposed contributions of $\overline{v^{2}}$ variances also demonstrate an energisation with $\overline{\beta}_{C,\theta}$ beyond the near-wall region, where most of the $\overline{v^{2}}$ energy is concentrated (Figures 3c,d). While $\overline{u^{2}}$ is strong in both the near-wall and outer regions, the majority of $\overline{v^{2}}$ energy is concentrated in the outer region of the TBL. This is attributed to the dominance of large-scale energy transfer through the pressure-strain mechanism in the outer layer, as demonstrated by Gungor et al. (2022) and Deshpande and Vinuesa (2024). Inner-scaled turbulence statistics in the lower-defect cases (Cases 1 and 2) show that accumulated PG and local disequilibration primarily affect the outer region, while the near-wall statistics remain largely unchanged. Since the $\overline{v^{2}}$ footprint is confined mainly to the outer layer, the entire wall-normal extent of $\overline{v^{2}}$ is sensitive to the effects of accumulated PG and local disequilibration.
Figure 4: Contours of the premultiplied spanwise power-spectral density of (a) streamwise Reynolds stress scaled with friction velocity $k_{z}\phi_{uu}/U_{\tau}^{2}$, and (b) wall normal Reynolds stress scaled with friction velocity $k_{z}\phi_{vv}/U_{\tau}^{2}$ corresponding to Case DISEQ as per Table 2. Figure also presents (c) Small scale contributions to $\overline{u^{2}}$ variances, (d) Large scale contributions to $\overline{u^{2}}$ variances, (e) Small scale contributions to $\overline{v^{2}}$ variances, (f) Large scale contributions to $\overline{v^{2}}$ variances. The dotted line corresponds to the $\lambda_{z,c}^{+}=300$ scale cutoff as per Deshpande et al. (2023). Colors and Line styles correspond to those in Tables 1 and 2.
@@
To further understand the effects of local disequilibration, particularly on the turbulent scales, we analyze the premultiplied spanwise power-spectral density scaled in inner units for Case DISEQ in Figure 4 (Table 2).
The impact of local disequilibration is most evident by observing the differential behavior of small- and large-scale energy contours (left and right of the dotted line marking $\lambda^{+}_{z,c}=300$ respectively). As shown in Figure 4(a), the $\overline{u^{2}}$ spectra for DNS22 exhibits a more energized inner peak than b1.4, indicating the rapid response of near-wall small scales to local PG changes. Meanwhile, the outer region of both $\overline{u^{2}}$ and $\overline{v^{2}}$ spectra (which includes both small and large scales), and the large scales in the inner region, reflect the delayed turbulence response associated with high disequilibration. The scale-decomposed variances shown in Figure 4 (bottom row) further support the spectral observations, with DNS22 $\overline{u^{2}_{SS}}^{+}$ variances (in Fig 4(c)) displaying near-wall energization but a delayed response in the outer region. These contrasting responses of the near-wall and outer regions are due to differences in their dominant turbulence scales; the near-wall region is dominated by physically small scales which have small time scales that respond quickly to any change in the PG conditions (Bobke et al., 2017). On the other hand, large scales that predominate in the outer region have a longer response time scale, leading to delayed adaptation. An exception, however, is the behaviour of small-scales in the outer region, which do not react as quickly to disequilibration as those in the near-wall region. Understanding the physical reasoning behind this warrants further investigation into their behavior, which is, however, beyond the present scope.
@@ 3.2 TBLs imposed with strong adverse pressure gradients
For Case 3, $Re_{\theta}$ and $\beta_{ZS}$ are matched, but the higher value of $\overline{\beta}_{ZS,\theta}$ in DNS22 suggests that it experiences a more pronounced historical effect of the APG than b1.4. Indeed, figure 5(a) shows the mean velocity defect increasing with $\overline{\beta}_{ZS,\theta}$, indicating a greater cumulative influence of the APG, consistent with our observations for $\overline{\beta}_{C,\theta}$ in Figures 2(top). Similarly, Figures 5(b) and 5(c) respectively show increasing outer peaks in the $\overline{u^{2}}$ and $\overline{v^{2}}$ variances with an increase in the $\overline{\beta}_{ZS,\theta}$ parameter. It is important to note here that the choice of an outer scaling-based PG parameter restricts our analysis solely to the outer region of the boundary layer, preventing reliable inferences about the inner region. We note that although the accumulated PG parameter $\overline{\beta}_{ZS,\theta}$ qualitatively aligns with the differences in the statistics, and is consistent with previously established trends of $\overline{\beta}_{C,\theta}$, the effectiveness of using $\overline{\beta}_{ZS,\theta}$ together with matched values of $Re_{\theta}$ and $\beta_{ZS}$ cannot be conclusively determined based on this case alone.
@@
Figure 6: Contours of the premultiplied spanwise power-spectral density of streamwise and wall normal Reynolds stresses scaled in outer units ($k_{z}\phi_{uu}/U_{e}^{2}$$k_{z}\phi_{vv}/U_{e}^{2}$) for Case 3 (a,b), Case 4(c,d), and Case 5(e,f). The dotted line corresponds to the $\lambda_{z,c}/\delta=0.5$ scale cutoff computed based on Deshpande et al. (2023). Colors and Line styles correspond to those in Tables 1 and 2.
@@
In terms of influence of local disequilibration, one can argue these effects are likely not to influence the statistics in Case 5 due to ${\rm d}{\beta}/{\rm d}{Re}$ values being of the same order of magnitude for the associated data sets (refer Table 2). For Cases 3 and 4, however, local disequilibration may play a more significant role, given that the ${\rm d}{\beta}/{\rm d}{Re}$ values differ by at least an order of magnitude between the two data sets being compared. Strong negative ${\rm d}{\beta_{ZS}}/{\rm d}{Re}_{\theta}$ for DNS22 (Case 3) and DNS16 (Case 4) imply a decreasing impact of the APG (Gungor et al., 2024b), which acts in opposition to the strong positive value of $\overline{\beta}_{ZS,\theta}$. The results for Cases 3 and 4 in Figure 4(a-f) indicate that the accumulated PG effect dominates, rather than the local disequilibration.
With the efficacy of $\overline{\beta}_{ZS,\theta}$ and $\overline{\beta}_{ZS,ZS}$ to quantify history effects demonstrated, next we leverage this new capability to understand the historical PG trends in power-spectral density and scale-decomposed contributions to velocity variances for these strong APG TBLs.
Figure 7: Small and large scale contributions to velocity variances in Cases 3 (top), 4 (middle), and 5 (bottom). Decomposition into small and large scales is done based on the $\lambda_{z,c}/\delta=0.5$ cutoff discussed in $\S$$2.$ All quantities are scaled with respect to outer units, $U_{e}$ and $\delta$. (a,e,i) Small scale contributions to $\overline{u^{2}}$ variances, (b,f,j) Large scale contributions to $\overline{u^{2}}$ variances, (c,g,k) Small scale contributions to $\overline{v^{2}}$ variances, (d,h,l) Large scale contributions to $\overline{v^{2}}$ variances. Colors and Line styles correspond to those in Tables 1 and 2.
Figure 6 presents premultiplied spanwise power-spectral density scaled in outer units for $\overline{u^{2}}$ and $\overline{v^{2}}$ in Cases 3, 4, and 5, complementing earlier inferences from velocity variances. The dotted line in Figure 6 marks the scale cutoff ($\lambda_{z,c}/\delta=0.5$, computed based on Deshpande et al. (2023)), to distinguish between small and large scales. Figure 6 for Cases 3, 4, and 5 shows similar trends of increasing outer-region energy with higher $\overline{\beta}_{ZS,\theta}$ and $\overline{\beta}_{ZS,ZS}$ values respectively, as observed in the low defect Case 2. In Cases 3 and 4, DNS22 and DNS16 displays outer-region small-scale energization and broader spectral spread, respectively, consistent with their higher $\overline{\beta}_{ZS,\theta}$, dominating the reduced APG effect caused by stronger negative disequilibration.
Case 5 reinforces this behavior - DNS16 exhibits increased energy at both ends of the spanwise wavelength spectrum for $\overline{u^{2}}$ and $\overline{v^{2}}$. Figure 6(e) shows a more energized and broader outer peak in $\overline{u^{2}}$ for DNS16, while Figure 6(f) captures a distinct outer peak in $\overline{v^{2}}$ that is absent in DNS22, both aligning with its greater $\overline{\beta}_{ZS,ZS}$ value.
@@
Figure 7 presents scale-decomposed contributions to $\overline{u^{2}}$ and $\overline{v^{2}}$ variances for Cases 3 to 5. We reiterate here that outer scaling limits our analysis of flow physics solely to the outer region. The plots reveal outer region energisation of both small- and large-scale contributions of $\overline{u^{2}}$ and $\overline{v^{2}}$ with increasing $\overline{\beta}_{ZS}$ (2 and 3), across all three high-defect Cases 3-5, consistent with Case 2 (Figure 2 (below)). Further, for Cases 3 and 4, the trend associated with the local disequilibration ($d{\beta}/d{Re}$; for both small and large-scales) is also consistent with that noted for Case 2 in the outer region.